OpenStudy (anonymous):

If x varies inversely with y and x = 8 when y = 6, find y when x = 10. a. y - 5/12 b. y = 4.8 c. y = 40/3 d. y = 7.5

OpenStudy (michele_laino):

Please note that you have to solve this equation: \[6 \cdot 8 = 10 \cdot y\] what is y?

Nnesha (nnesha):

do you know the equation that u have to use for inverse variation ??

Nnesha (nnesha):

\[\huge\rm y = \frac{ k }{ x }\] this is equation that u have to use now replace x and y by their values and solve for k

OpenStudy (anonymous):

I think it is D but I am not sure

Nnesha (nnesha):

nope

Nnesha (nnesha):

 use first x and y values to solve for k